OpenStudy (anonymous):

find the zeros of f(x)= x^2-4x-12

hartnn (hartnn):

to get zeros of f(x), equate f(x) =0 so can you solve this quadratic , x^2-4x-12=0 ??

OpenStudy (anonymous):

so i would add -12to zero

OpenStudy (anonymous):

then dived by -4?

OpenStudy (campbell_st):

no, find the factors of -12 that add to -4, the larger is negative and the smaller is positive.

hartnn (hartnn):

2 numbers whose sum is -4 and product is -12 are.... ?

OpenStudy (anonymous):

-6 and 2

hartnn (hartnn):

correct! so write -4x as -6x+2x x^2-6x+2x -12 =0 factor out x from 1st 2 terms and +2 from last 2 terms, what u gt ?

OpenStudy (anonymous):

This is a quadratic. You cannot solve a quadratic by simple manipulation (adding to both sides, multiplying, etc) Instead, you have a few options: Factoring Quadratic formula Completing the square

hartnn (hartnn):

we are doing factoring method

OpenStudy (anonymous):

You'll probably end up learning all of those methods in your class. Which have you studied so far? I would imagine factoring is how your teacher expects you to be doing quadratics so far. Factoring is much simpler to do, but it doesn't always work the way the other two methods always work. Sometimes the quadratic will be unfactorable.

OpenStudy (anonymous):

i sorta remember the 6x + 2x part bur not really. i don't understand how to factor it beyond that

hartnn (hartnn):

ok, take \(\large x^2-6x\) part it can be written as \(\large x \times x-6\times x=x \times (x-6)\) got this ? this is factoring x out from 1st 2 terms ...

hartnn (hartnn):

nowcan u fatcor +2 from +2x-12 ?

OpenStudy (anonymous):

10-x?

hartnn (hartnn):

10? where does 10 come from ?

OpenStudy (anonymous):

i don't know? nevermind?

hartnn (hartnn):

okk.. \(\large 2x-12 = 2\times x-2\times 6 = 2\times (x-6)\) its just opposite of distributing...u getting this ?

OpenStudy (anonymous):

yeah i see what you are saying

OpenStudy (anonymous):

Okay, so think of factoring this way: You're setting up two parenthesis, and then you are trying to fill in those parenthesis so that when you multiply them together, you would get what you started with. Let me show you an example. \(\Large x^2 + 7x + 10\) \(\Large(\text{_ }\pm\text{ _})(\text{_ }\pm\text{ _})\)

OpenStudy (anonymous):

Now, think about how those parenthesis would get multiplied. Have you practiced FOIL?

hartnn (hartnn):

good, so we are here now \(x(x-6)+2(x-6)\) by exactly same logic, factor out x-6, try ?

OpenStudy (anonymous):

yes when i was in algerbra 2 two years ago

OpenStudy (anonymous):

Haha okay so do you remember FOILing? Firsts, then Outer and Inner, and then Last.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, so the Firsts part of the multiplying will give you the \(x^2\) part, so in the firsts place of the parenthesis, we put something that will multiply to be \(x^2\).

OpenStudy (anonymous):

\((\text{x }\pm\text{ _})(\text{x }\pm\text{ _})\)

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

no? sorry

OpenStudy (anonymous):

Being good at factoring is really all about understanding FOIL well, so as much work as it might seem, you should go back and make sure you're good at that. Maybe I need to reexplain it to you.

OpenStudy (anonymous):

oh wait, zeros are -2 and +6

OpenStudy (anonymous):

Why do you say that? How did you get those numbers?

OpenStudy (anonymous):

i factored (x+2)(x-6)

OpenStudy (anonymous):

Okay, cool. You factored correctly =) To check that it's factored correctly, you would foil that. \((x+2)(x-6)\) Firsts gives you x*x = \(x^2\) Outer gives you x*-6 = -6x Inner gives you 2*x = 2x Lasts gives you 2*-6 = -12 So \(x^2 -6x +2x - 12 = x^2 -4x - 12\) Which means it's factored correctly.

OpenStudy (anonymous):

okay awesome thanks

OpenStudy (anonymous):

Now factoring doesn't get you the answers immediately, but it gets you super close!

OpenStudy (anonymous):

Now that we've factored, the equation we have is \(\large (x+2)(x-6) = 0\) think about that equation. I have two things, and I'm multiplying them together, and they equal zero.

OpenStudy (anonymous):

What could those things be? How can I have two things, and multiply them together and get 0?

OpenStudy (anonymous):

Hello? Thoughts?

OpenStudy (anonymous):

Can we finish this question? It pains me to see you get 95% there and then not cross the finish line.

OpenStudy (anonymous):

@sarahjones667

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Okay, so I multiply two things together, and the result is 0. Can you tell me anything about the two things?

OpenStudy (anonymous):

they are 0?

OpenStudy (anonymous):

Almost correct. They don't have to both be 0, but at least one of them is.

OpenStudy (anonymous):

one is 0 one is 1

OpenStudy (anonymous):

Does that make sense? So if I have (x+2)(x-6) = 0 What I can decide is, either (x+2) = 0 or (x-6) = 0 otherwise, multiplying them together wouldn't give me 0.

OpenStudy (anonymous):

Does that much make sense?

OpenStudy (anonymous):

yeah it does

OpenStudy (anonymous):

Cool, so all we have to do finally is look at those two equations and figure out the x value for either one. x+2 = 0 what is x?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

That's your first zero =)

OpenStudy (anonymous):

And your second zero?

OpenStudy (anonymous):

+6

OpenStudy (anonymous):

Good work =) So cool. In the future, factor just like that, but remember that you don't get your answer until you figure out what x value would make those factors equal 0.

OpenStudy (anonymous):

okay thank ya thank ya

OpenStudy (anonymous):

My pleasure =)